### v1

# Tangling-Untangling Cycle for Efficient Learning

Abstract
The conventional wisdom of manifold learning is based on nonlinear dimensionality reduction techniques such as IsoMAP and locally linear embedding (LLE). We challenge this paradigm by exploiting the blessing of dimensionality. Our intuition is simple: it is easier to untangle a low-dimensional manifold in a higher-dimensional space due to its vastness, as guaranteed by Whitney embedding theorem. A new insight brought by this work is to introduce class labels as the context variables in the lifted higher-dimensional space (so supervised learning becomes unsupervised learning). We rigorously show that manifold untangling leads to linearly separable classifiers in the lifted space. To correct the inevitable overfitting, we consider the dual process of manifold untangling - tangling or aliasing- which is important for generalization. Using context as the bonding element, we construct a pair of manifold untangling and tangling operators, known as tangling-untangling cycle (TUC). Untangling operator maps context-independent representations (CIR) in low-dimensional space to context-dependent representations (CDR) in high-dimensional space by inducing context as hidden variables. The tangling operator maps CDR back to CIR by a simple integral transformation for invariance and generalization. We also present the hierarchical extensions of TUC based on the Cartesian product and the fractal geometry. Despite the conceptual simplicity, TUC admits a biologically plausible and energy-efficient implementation based on the time-locking behavior of polychronization neural groups (PNG) and sleep-wake cycle (SWC). The TUC-based theory applies to the computational modeling of various cognitive functions by hippocampal-neocortical systems.
I Introduction
A key challenge in machine learning (ML) is balancing specification and generalization (a.k.a. over-fitting and under-fitting). We want a model that generalizes well (but not so simplistic that it underfits) and is well-specified to capture the essential patterns in the training data (but not so complex that it overfits). This subtle balance is often achieved through statistical techniques like cross-validation, regularization, and choosing the right model complexity. From a geometric perspective, the challenge is intrinsically related to the manifold structure of the data acquired from the physical world. To discover and exploit such manifold structure, nonlinearity is commonly believed to be essential, as witnessed by kernel trick in support vector machine [48] and nonlinear dimensionality reduction (e.g., IsoMAP [45] and locally linear embedding [40]). However, nonlinearity is often at odds with rigorous proof and the simplicity criterion in mathematical and scientific research.
In this paper, we draw inspiration from the discovery of place cells [34] and its related aliasing problem [29, 50]. Nature has discovered an elegant solution to efficient learning through the evolution of mammalian brains, but its secret has remained elusive. We shed some light on this greatest puzzle by revisiting the role of specification and generalization from a manifold untangling/tangling perspective [27]. Whitney’s embedding theorem in topology [31] has shown that it is easier to untangle a manifold in a high dimensional space. Such intuition inspires us to introduce context dependency as the clue for manifold untangling. Using class labels C as contextual variables, we can show an arbitrary manifold $P(\textbf{X})$ can be untangled (admitting linear dichotomy [6]) in the lifted space $P(\textbf{X},\textbf{C})$. The opposite direction (tangling operator) is a simple integral transform that collapses/aliases all variables sharing the same context. By associating tangling/untangling with context-independent representation (CIR) and context-dependent representation (CDR), we rigorously show how to strike an optimal tradeoff between specificity and generalization by a pair of tangling and untangling operators, which we call tangling-untangling cycle (TUC).
Using TUC as the building block, we can construct more sophisticated learning algorithms based on Cartesian products and fractal geometry [28]. It can be shown that the Cartesian product of TUC is still a TUC, which supports the multiscale extension from low-dimensional to high-dimensional space. To mirror the nested structure of data acquired from the physical world, fractal-inspired extension is constructed where the inter-scale context dependency is induced by an index variable (a.k.a. positioning trick [26]). The efficiency of the proposed learning algorithms is supported by their biological implementations based on polychronization neural groups (PNG) [24]. The implementation of TUC by sleep-wake cycle [13] finishes our journey at its starting point - the discovery of place cells in the hippocampus [34]. The presence and absence of external stimuli (i.e., default state vs. perturbation state) offer an energy-efficient and biologically plausible implementation of untangling and tangling operators in wake and sleep modes, respectively. Overall, the main contributions of this paper can be summarized as follows:
  • A geometric insight about the blessing and curse of dimensionality leads to a new understanding of specification and generalization in ML, which nicely fits the emerging direct-fit-to-nature paradigm.
  • A construction of manifold tangling-untangling cycle (TUC) for efficiency learning. Our construction represents a novel direct-fit-to-nature approach to unsupervised learning. Tangling and untangling operators are implemented by context dependency and integral transform, respectively.
  • A biological connection of TUC with sleep-wake cycle (SWC) and its related cognitive functions. Untangling transforms CIR to CDR for specification in the wake mode; and untangling maps CDR back to CIR for generalization in the sleep mode.
  • An application of TUC theory to model sensorimotor and social interactions. We use biological findings in various studies of cognitive functions as experimental support for the new theoretical framework.
II Geometric Intuition
Figure 1: Kernel trick achieves linear separability by mapping input data to a higher-dimensional feature space. $\phi:(x,y)\rightarrow(x,y,c)$, where $c=0,1$ denote red and blue classes, is a simple lifting idea that can be generalizable to a data manifold in arbitrary dimensions.
In graph theory, there is a well-known educational puzzle named “Three Utilities Problem”. This puzzle can be formalized as a problem in topological graph theory by asking whether the complete bipartite graph $K_{3,3}$, has a graph embedding in the plane. The impossibility of the puzzle corresponds to the fact that $K_{3,3}$ is not a planar graph. However, such a puzzle can be easily solved if the planarity constraint is removed. This is a simple example illustrating the blessing of dimensionality to the problem of manifold untangling. In topology, Whitney Embedding Theorem [31] states that any smooth manifold can be embedded into a Euclidean space of double its dimension. Based on this classical result, we can prove the following lemma:
Lemma (Linear separability in high-dimensional space)
Given two tangled manifolds $A,B\in R^{n}$, one can always linearly separate them in a high-dimensional space.
Proof. Using Whitney’s embedding theorem, we can also embed $A,B$ to a Euclidean space in $\tilde{A},\tilde{B}\in R^{2n}$. Then it is trivial to linearly separate $\tilde{A},\tilde{B}$ in $R^{4n}$.
In machine learning (ML), the kernel trick used by support vector machine [42] represents a good example for illustrating the blessing of dimensionality. As shown in Fig. 1, linearly non-separable data sets can become untangled (linearly separable) after being mapped to a high dimensional feature space. The construction or learning of features has been at the foundation of all existing models, including both biological and artificial neural networks. Recent direct-fit models suggest that we can solve the manifold untangling problem without feature engineering [20]. This view is also consistent with intelligence without representation [3] and supported by the discovery of concept cells [38, 37]. As summarized in [9], identity-related feature information of an object manifold can be simultaneously conveyed by the gnostic units in the IT cortex along with contextual information (e.g., position, pose, and scale), implying a high-dimensional latent space.
III Context Dependency and Integral Transform for Untangling and Tangling
Our intuition is based on observing the duality between specification (overfitting) and generalization (underfitting) in learning. An alternative approach to over/under-fitting is direct-fit-to-nature [20], which greatly inspired our approach in this work. Geometrically, we interpret overfitting/underfitting as manifold untangling and tangling, as shown in Fig. 2. Assuming sensory raw inputs are data belonging to a low-dimensional manifold embedded in a high-dimensional space, the objective of manifold untangling is to overfit the data for specification. This objective can be facilitated by embedding the manifold into a high-dimensional space, as Whitney’s embedding theorem suggests. Therefore, manifold untangling maps sensory raw input to a high-dimensional latent space where the fitting task becomes trivial (e.g., by linear dichotomy [6]). The dual process of manifold tangling is to underfit the data for generalization. A common example is aliasing [29], which reduces the dimensionality by collapsing the variables with a common state. We suggest an integral transform as a more general strategy than aliasing for generalization. In this section, we will study the pair of manifold untangling and tangling separately.
Figure 2: The tangling-untangling cycle (TUC) for solving overfitting/underfitting problems in efficient ML. Note its biological plausibility is consistent with sleep-wake cycle (SWC).
III-A Manifold Untangling: Dimensionality Expansion via Context Dependency
The intuition behind context dependency can be easily explained by the toy example of binary classification (e.g., cats vs. dogs). Suppose we collect sensory raw data associated with two classes $\textbf{X}_{0},\textbf{X}_{1}\in R^{n}$ corresponding to classes 0 and 1, respectively. The raw data in $R^{n}$ is tangled (e.g., blue vs. red dots as shown in Fig. 1) but they can be easily untangled by lifting the dimensionality by one - i.e., $(\textbf{X}_{c},c)\in R^{n+1}$ where $c=0,1$ (a linear classifier $c=1/2$ is sufficient). Such an idea of dimensionality expansion can be better articulated as a strategy of context dependency where the label is explicitly coupled with the training data as a context variable. Note that the introduction of label-based context dependency transforms supervised learning to unsupervised learning [2] by direct-fit-to-nature [20]. We can generalize the strategy of context dependency for high-dimensional labels/context as follows.
Proposition 1 (Specification). Context Dependency for Manifold Untangling.
Given sensory raw data $\textbf{X}\in R^{n}$ and its associated label $\textbf{C}\in R^{k}$, we can untangle the tangled data distribution $P(\textbf{X})$ by introducing context dependency as a joint PDF $P(\textbf{X},\textbf{C})\in R^{n+k}$.
The proof of proposition 1 is a straightforward extension of the strategy used in binary classification. By considering the discrete representation of label or context C, linear separability is guaranteed in the expanded higher-dimensional space.
[…]
III-B Manifold Tangling for Generalization: Dimensionality Reduction via Integral Transform
The dual operation of manifold tangling can be understood as the extension of aliasing [29]. The basic idea behind aliasing is to provide useful invariants by representing as equivalent states in which the same action/class is needed. Aliasing is good for generalization but has the risk of over-generalization. Such risk can be well managed by applying aliasing or its extended tangling operator to the untangled data in the high-dimensional space. The trick is to selectively remove or collapse hidden states where the context dependency dictates the criterion of selectivity. Using contextual information C as the guardrail, we are guaranteed a faithful generalization without the risk of over-generalization.
Proposition 2 (Generalization). Integral Transform for Manifold Tangling.
[…]
The pairing of manifold untangling and tangling operators forms a complete (TUC) cycle, as shown in Fig. 2. The connection of TUC with direct-fit models [20] is appealing - the objective of learning is not human-interpretable rules or world representations (i.e., “intelligence without representation” [3]). Instead, evolution has invented a localized computation for embodied cognition to interpolate over task-relevant manifolds (we will elaborate on this issue later when studying sensorimotor interaction). Manifold untangling offers a constructive solution to obtain over-parameterized models that are simple and parsimonious. More importantly, the connection of TUC with the evolutionary process is so powerful that we strongly believe it offers an important clue to unlock hidden secrets in developmental and ecological psychology. In summary, TUC forms the building block for our solution to efficient learning problems in general. Before we study the application of this new theory into neuroscience, we report several hierarchical extensions of TUC into chains and fundamental group of torus [21] to enrich our weaponry of constructing over-parameterized direct-fit models.
IV Hierarchical Extension and Connection with Fractal Geometry
First, we can spatially extend TUC with Cartesian products to form a fundamental group of the circles [21]. Such a dimensional increase is similar to the product PDF from the two marginals. Note that after manifold untangling, factorized linear dichotomies are conceptually similar to naive Bayes [39] (simple and parsimonious low-dimensional classifiers). Therefore, the Cartesian product preserves the independence of factorized distributions in the latent untangled space. In this manner, direct-fit models can be arbitrarily scaled to a high-dimensional space without suffering from the curse of dimensionality. Formally, we have the following result:
a)                                                      b)
Figure 3: Hierarchical extension of TUC: a) spatial extension by Cartesian product of TUC, which forms the fundamental group of the circles; b) temporal extension by alternating between untangling and tangling in time domain.
Proposition 3. (Cartesian Product of TUC forms the Fundamental Group of the Circle).
[…]
By recursively applying Cartesian products of TUC, we can construct a topological structure with increasing dimensionalities. Fig. 3a illustrates a toy example for such an extension of increasing dimensionalities (from a pair of circles to a torus). Analogous to the extension of circles to the fundamental group of circles, we can construct the fundamental group of TUC as $\textbf{S}^{1}\times...\times\textbf{S}^{1}$.
An alternative hierarchical extension of TUC is based on the temporal depth [35] where the process of untangling and tangling occurs in an alternating manner. Temporal depth or hierarchy is formed based on allostasis [35], as shown in Fig 3b. This construction will be useful for characterizing the dynamics of TUC, such as error feedback and temporal stability, especially in the presence of heterogeneous interactions (e.g., sensory interactions with motor). The alternating between tangling and untangling operators is also analogous to the alternating between sleep and wake modes in sleep-wake cycles, as we will elaborate in the next section.
Proposition 4. (Allostatic concatenation of TUCs forms temporal hierarchy.
Temporally hierarchical extension of TUCs can support continual learning by alternating between the processes of untangling and tangling in the time domain.
A more subtle construction of hierarchical TUC is based on fractal geometry [28]. We must extend the strategy context dependency across different scales to support fractal TUC. This can be implemented by the positioning trick [26], which was inspired by Hippocampal indexing theory [46]. The indexing variables $I$ induce inter-scale context dependency - e.g., in structure-mapping theory for analogy [16], higher-order relations are modeled based on the relation between objects instead of attributes of objects (so-called systematicity principle). The implementation of such principle requires the context to be the object X instead of class label C, which requires a fractal design such as the six-layered organization of cortical column. Such hierarchical organization of structures into TUCs might explain the magical number of six (e.g., six degrees of separation) or seven [30].
Figure 4: Illustration of sleep-wake cycle (SWC): untangling occurs during the wake time (perturbation state) and tangling occurs during the sleep time (default mode).
V Biological Connections with PNG and SWC
The biological implementation of the untangling operator is based on the connection of the polychronization neural groups (PNG) theory in [24] with the locally subspace untangling hypothesis in [9]. A direct consequence of direct-fit-to-nature is over-parameterization [20] to increase the predictive power without explicitly modeling the underlying generative structure of the world (i.e., intelligence without representation [3]). PNG offers a plausible computational model for supporting such mindless yet powerful generalizations. The number of PNG [24] can exponentially increase as more neurons are involved. Intuitively, one can argue that a high-level PNG with more neurons will have more representation power than a low-level PNG, which makes it a promising candidate for implementing the direct-fit-to-nature model [20].
We first extend Hebbian postulate [22] to make it work for PNG as follows - “A low-level PNG that $drives$ a high-level PNG to fire tends to wire together” [11]. Driving is hypothesized as a mechanism for generalizing cloning [50] to achieve specificity (e.g., unidirectional place cells [4] and context-dependent concept cells). The opposite direction from high-level to low-level generalizes aliasing [29] to achieve invariance or generalization (e.g., omnidirectional place cells and context-independent concept cells [38, 37]). Due to dimensionality expansion (the consequence of recursive untangling), the support of CDR would increase exponentially. PNG-based implementation can handle this combinatorial explosion because of its enormous capacity [24].
[…]
a)                                                      b)
Figure 5: Context dependency matters to both visual perception and motor control. a) PNG-based untangling for visual perception; b) PNG-based untangling for motor control.
The combination of PNG with SWC [13] allows us to study cognitive functions from a novel TUC perspective. Using TUC as the vocabulary, we attempt to construct a computational model for SWC as the first step toward a more comprehensive neural syntax [4]. As application examples, we consider three tasks here: spatial navigation, object recognition, and motor control.
Spatial Navigation. Navigation represents an easier learning problem that evolution solved long before the vision problem. Our understanding about this task has converged to a theory based on cognitive maps [47, 34, 50]. Cognitive maps organize knowledge to support flexible and adaptive behavior. How do mammalian brains encode 2D physical space information into 1D spike sequence? This question is at the core of understanding the relationship between space and time in its simplest form. The representation of spatial information (Euclidean or non-Euclidean) plays an essential role in the building of cognitive maps. A graph-based representation such as clone-structured causal graphs (CSCG) [17] has been widely adopted for encoding the physical space into action-related state space. During the construction of the CSCG model, the aliasing problem is solved by the cloning trick, which can be interpreted as a degenerated untangling operator.
Untangling during the wake mode (perturbation state) is demonstrated by place cells with unidirectional place fields. The 1D travel sequence dictates the spatio-temporal context, which helps resolve the ambiguity in space [4]. Tangling during the sleep mode (default state) is demonstrated by omnidirectional place cells [34] that do not require a temporal context. Memory consolidation [44, 49] implements the aliasing or tangling by binding all PNG which shared the same context C together. The collapse of cloned variables occurs at the hippocampus where the encoding of spatio-temporal context (index to the local map or position on the global map) is handled by CA3-CA1 collateral system (pp. 326 [4]). During the consolidation of semantic memory (collection of CIR) from episodic memory (collection of CDR), the PNG-based time-locking mechanism eliminates the spatio-temporal context to support context-independent inference for generalization.
[…]
Motor control. The duality between visual perception and motor control was observed in [36]. Based on this analogy, one can understand the untangling problem in the motor cortex similarly to that in the sensory cortex [51], as shown in Fig. 5b. The high-tangling state consists of options that are context-independent; while the low-tangling state is the collection of actions that are context-dependent. The importance of motor untangling to muscle-like actions/commands has been demonstrated in [41]. Sleep mode plays an important role in motor learning similarly to visual learning. During sleep, particularly during rapid eye movement (REM) and deep stages of non-REM sleep, the brain consolidates motor skills learned during the day. This process is essential for motor learning, helping to improve skills and refine motor tasks over time. Empirical evidence supporting this line of theoretical reasoning can be found in [7].
VI Computational Modeling of Sensorimotor and Social Interactions
TUC-based local computation for direct-fit inspired us to propose a new conceptual framework called “localized embodied cognition” with the locality principle in mind from problem formulation to solution algorithms. We still use the world “as its own model” [3], but the way local embodied cognition never attempts to solve a global problem locally. An organism needs to couple sensory with motor systems to achieve various goals but never stores knowledge about the global environment. Instead, the bodily interactions with the environment (situatedness) are persistently maintained by the neocortex (a single book)’s interaction with the hippocampus (a large library). The cognitive map in the hippocampus offers a situational context to interpret the sensorimotor interaction occurring in the neocortex. In this section, we study sensorimotor interaction and hippocampal-neocortical interaction using the newly-constructed neural syntax.
Figure 6: Sensorimotor cycle (SMC) consists of sensory untangling-tangling pair and motor untangling-tangling pair for predictive processing. Note that sensory and motor context-dependent representations (CDR) are task-relevant latent variables (the connection of chain); sensory and motor context-independent representation (CIR) are the interface with the external environment.
Neocortex as a Sensorimotor Machine.
One of the biggest mysteries in studying cortical columns is that they are structures without a function [23]. Instead, several canonical microcircuits such as Douglas-Martin model [10] and Haeusler-Maass model [18] have been studied for predictive coding in the literature. These models have revealed a remarkable correspondence between the microcircuitry of the cortical column and the connectivity required by predictive coding. Cortical columns do not have a function because they work like swiss-army knivies. By connecting with different sensory and motor areas, they offer a universal processing algorithm [32]. TUC-based modeling allows us to gain a deeper understanding of sensorimotor interaction at the system level without the distraction from lower-level implementations.
As shown in Fig. 6, the coupling of sensory and motor untangling-tangling pairs naturally forms a bigger cycle for the next level along the hierarchy of system neuroscience - modeling sensorimotor interaction by a pair of inverse and forward models [25]. Since more is different [1], the sensory motor cycle (SMC) is likely to be guided by a varying organizational principle, as initially demonstrated by [43]. Based on the duality between sensory and motor cortex [36], sensory CDR (after perceptual untangling) has to be linked with motor CDR (after motor untangling). They are task-relevant latent variables in the network dynamics of SMC [43]. We come up with the following conjecture but leave computational modeling for the mechanism of inverse and forward models in SMC and its relationship to SWC for a future study.
Conjecture (Context-dependent manifold untangling) The task of sensorimotor learning can be facilitated by simultaneously untangling the neural manifold of sensory and motor variables in the right positioning of latent space, where the position is specified by the index from the hippocampus as in Hippocampal indexing theory [46].
SMC-based Social Interaction
. Carrying the “more is different” principle one step further, we can envision a bigger cycle connecting a pair of SMCs. When one person becomes another’s environment, the communication between two SMCs forms another bigger cycle known as non-verbal communication. The communication bandwidth of such non-verbal communication is limited by the locality of our sensory organs. Even though within each neocortex, the hierarchical organization maximizes the information flow inward, social communication (flowing outward) has to wait for evolution to improve its efficiency.
Figure 7: The cycle connecting speech production by motor cortex and language comprehension by auditory cortex has a dual structure to SMC because social communication between two people is connected by CIR (tangled states).
The fundamental role played by natural language is to facilitate communication among people - from non-verbal to verbal. Fig. 7 shows a modification of SMC to abstract the verbal communication between two people A and B, where motor and auditory cortex are responsible for speech generation and language comprehension, respectively. Note that unlike the original SMC in Fig. 6, the interface between two TUCs is in the tangled state (CIR). This difference has a fundamental impact on our internal (self-conscious) and external (social) cognitive function. For internal communication between sensory and motor cortex, the interface (inverse and forward in Fig. 6) connects two untangled states and there is no bottleneck in terms of information flow. For external (social) communication between two different person, the interface (inverse and forward in Fig. 7) connects two tangled states and faces an information flow bottleneck.
According to the simulation theory of mind-reading [15], mirror neurons can help us understand and empathize with the emotions and intentions behind others’ non-verbal cues. Modeling the new cycle in Fig. 7for understanding the social nature of language might offer new insight to the neural exploitation hypothesis [14]. According to this hypothesis, social cognition and language are both connected with action through the support of mirror neurons. However, the difference between non-verbal and verbal communication seems to be determined by their organizational principles. Non-verbal communication such as sign language does not admit hierarchical organization; while verbal communication such as Chomsky’s hierarchy [5] in natural language offers an evolutionary solution to maximize the information flow within the new cycle.
VII Conclusions and Future Works
In this paper, we have constructed a new TUC-based theory to solve the long-standing learning problem. Exploiting the blessing of dimensionality by inducing context dependency allows us to untangle any data manifold so linear classifiers will be sufficient. Then the tangling operator based on integral transform collapses the induced context variables for generalization. Such mapping between tangled CIR and untangled CDR can be extended into hierarchical representations by Cartesian product and fractal geometry. The computational efficiency of TUC-based learning algorithms is supported by the biological implementation by SWC. Sleep and wake modes in SWC elegantly correspond to tangling and untangling operators due to the absence of sensory inputs during the sleep mode. SWC allows us to construct a neural syntax for an improved understanding of cognitive functions associated with sensorimotor interaction and social interaction.

### v2

@@
-Tangling-Untangling Cycle for Efficient Learning
+On Computational Modeling of Sleep-Wake Cycle
 Abstract
-The conventional wisdom of manifold learning is based on nonlinear dimensionality reduction techniques such as IsoMAP and locally linear embedding (LLE).
-We challenge this paradigm by exploiting the blessing of dimensionality.
-Our intuition is simple: it is easier to untangle a low-dimensional manifold in a higher-dimensional space due to its vastness, as guaranteed by Whitney embedding theorem.
-A new insight brought by this work is to introduce class labels as the context variables in the lifted higher-dimensional space (so supervised learning becomes unsupervised learning).
-We rigorously show that manifold untangling leads to linearly separable classifiers in the lifted space.
-To correct the inevitable overfitting, we consider the dual process of manifold untangling - tangling or aliasing- which is important for generalization.
-Using context as the bonding element, we construct a pair of manifold untangling and tangling operators, known as tangling-untangling cycle (TUC).
-Untangling operator maps context-independent representations (CIR) in low-dimensional space to context-dependent representations (CDR) in high-dimensional space by inducing context as hidden variables.
-The tangling operator maps CDR back to CIR by a simple integral transformation for invariance and generalization.
-We also present the hierarchical extensions of TUC based on the Cartesian product and the fractal geometry.
-Despite the conceptual simplicity, TUC admits a biologically plausible and energy-efficient implementation based on the time-locking behavior of polychronization neural groups (PNG) and sleep-wake cycle (SWC).
-The TUC-based theory applies to the computational modeling of various cognitive functions by hippocampal-neocortical systems.
-I Introduction
-A key challenge in machine learning (ML) is balancing specification and generalization (a.k.a.
-over-fitting and under-fitting).
-We want a model that generalizes well (but not so simplistic that it underfits) and is well-specified to capture the essential patterns in the training data (but not so complex that it overfits).
-This subtle balance is often achieved through statistical techniques like cross-validation, regularization, and choosing the right model complexity.
-From a geometric perspective, the challenge is intrinsically related to the manifold structure of the data acquired from the physical world.
-To discover and exploit such manifold structure, nonlinearity is commonly believed to be essential, as witnessed by kernel trick in support vector machine [48] and nonlinear dimensionality reduction (e.g., IsoMAP [45] and locally linear embedding [40]).
-However, nonlinearity is often at odds with rigorous proof and the simplicity criterion in mathematical and scientific research.
-In this paper, we draw inspiration from the discovery of place cells [34] and its related aliasing problem [29, 50].
-Nature has discovered an elegant solution to efficient learning through the evolution of mammalian brains, but its secret has remained elusive.
-We shed some light on this greatest puzzle by revisiting the role of specification and generalization from a manifold untangling/tangling perspective [27].
-Whitney’s embedding theorem in topology [31] has shown that it is easier to untangle a manifold in a high dimensional space.
-Such intuition inspires us to introduce context dependency as the clue for manifold untangling.
-Using class labels C as contextual variables, we can show an arbitrary manifold $P(\textbf{X})$ can be untangled (admitting linear dichotomy [6]) in the lifted space $P(\textbf{X},\textbf{C})$.
-The opposite direction (tangling operator) is a simple integral transform that collapses/aliases all variables sharing the same context.
-By associating tangling/untangling with context-independent representation (CIR) and context-dependent representation (CDR), we rigorously show how to strike an optimal tradeoff between specificity and generalization by a pair of tangling and untangling operators, which we call tangling-untangling cycle (TUC).
-Using TUC as the building block, we can construct more sophisticated learning algorithms based on Cartesian products and fractal geometry [28].
-It can be shown that the Cartesian product of TUC is still a TUC, which supports the multiscale extension from low-dimensional to high-dimensional space.
-To mirror the nested structure of data acquired from the physical world, fractal-inspired extension is constructed where the inter-scale context dependency is induced by an index variable (a.k.a.
-positioning trick [26]).
-The efficiency of the proposed learning algorithms is supported by their biological implementations based on polychronization neural groups (PNG) [24].
-The implementation of TUC by sleep-wake cycle [13] finishes our journey at its starting point - the discovery of place cells in the hippocampus [34].
-The presence and absence of external stimuli (i.e., default state vs.
-perturbation state) offer an energy-efficient and biologically plausible implementation of untangling and tangling operators in wake and sleep modes, respectively.
+Why do mammals need to sleep?
+Neuroscience treats sleep and wake as default and perturbation modes of the brain.
+It is hypothesized that the brain self-organizes neural activities without environmental inputs.
+This paper presents a new computational model of the sleep-wake cycle (SWC) for learning and memory.
+During the sleep mode, the memory consolidation by the thalamocortical system is abstracted by a disentangling operator that maps context-dependent representations (CDR) to context-independent representations (CIR) for generalization.
+Such a disentangling operator can be mathematically formalized by an integral transform that integrates the context variable from CDR.
+During the wake mode, the memory formation by the hippocampal-neocortical system is abstracted by an entangling operator from CIR to CDR where the context is introduced by physical motion.
+When designed as inductive bias, entangled CDR linearizes the problem of unsupervised learning for sensory memory by direct-fit.
+The concatenation of disentangling and entangling operators forms a disentangling-entangling cycle (DEC) as the building block for sensorimotor learning.
+We also discuss the relationship of DEC and SWC to the perception-action cycle (PAC) for internal model learning and perceptual control theory for the ecological origin of natural languages.
+1 Introduction
+We spend one-third of our life in sleep [9]; infants often sleep over half of the time every day.
+Why?
+As the brain’s default mode, sleep is believed to play an important role in the health and functioning of mammalian brains, encompassing several vital functions.
+Sleep deprivation often harms cognitive functions, including learning and memory.
+It is well known that memory consolidation, a process in which episodic memory becomes semantic memory, occurs during sleep [74].
+Rapid eye movement (REM) sleep is believed to be important for consolidating procedural memories, those related to skills and tasks performed automatically without conscious awareness [69].
+However, little is known about the mechanism of sleep, not to mention the computational modeling of the sleep-wake cycle (SWC) in the open literature.
+In this paper, we fill this gap by drawing inspiration from the discovery of place cells [55] and its related aliasing problem [46, 75].
+The aliasing problem, arising from the locality constraint of the receptive field, is solved by memory consolidation during sleep.
+Context-dependent unidirectional place cells in episodic memory are transformed into context-independent omnidirectional place cells in semantic memory [9].
+Such a procedure can be intuitively interpreted as a manifold disentanglement or dimensionality reduction for generalization.
+The example of place cells for spatial navigation inspires us to generalize the role of memory consolidation into other cognitive tasks, such as perception and action.
+More specifically, we introduce context [^1] Context is defined as a set of conditions in which an input is uniquely coupled to an output [9].
+dependency as design bias [25] or inductive bias [3] for entanglement (associated with specification or overfitting).
+Figure 1: Illustration of sleep-wake cycle (SWC): entanglement/tangling occurs during the wake time (perturbation state) and disentanglement/untangling occurs during the sleep time (default mode).
+Using motion as contextual variables, we can show an arbitrary manifold $P(\textbf{X})$ can be entangled during the wake mode into the lifted space $P(\textbf{X},\textbf{C})$ as a linearization strategy for classification, which is closely related to the recently proposed direct-fit strategy [30].
+The opposite direction (disentangling operator) is abstracted by a simple integral transform that collapses/aliases all variables, sharing the same context variables for generalization.
+By associating entangling/disentangling with context-independent representation (CIR) and context-dependent representation (CDR), we rigorously show how the sleep-wake cycle (SWC) [21] might computationally implement a disentanglement-entanglement cycle (DEC) as shown in Fig.
+1.
+In wake mode, the sensory cortex encodes the external environment $P(\textbf{Y})$ in CDR $P(\textbf{X},\textbf{C})$ in the latent space for specification during bottom-up processing; the projection from CDR $P(\textbf{X},\textbf{C})$ to CIR $P(\textbf{X})$ characterizes the recurrent top-down processing based on error feedback.
+When external environment Y is absent in the sleep mode, memory consolidation transforms CDR $P(\textbf{X},\textbf{C})$ into CIR $P(\textbf{X})$ by virtual “replay” for generalization.
+Using DEC as the building block, we can construct more sophisticated sensorimotor learning algorithms for computational modeling of hippocampal-neocortical systems [8, 41].
+In the framework of sensorimotor learning in the latent space [67], we present a computational model of the perception-action cycle (PAC) [22] based on the DECs of sensory and motor cortex, respectively.
+It is hypothesized that context dependency facilitates the task of sensorimotor learning in the latent space by linearizing error feedback in a task-optimal manner [54].
+This hypothesis is supported by the duality between visual perception and motor control [57].
+Similar to untangling in visual perception [16], disentanglement of actions into options has also been observed in motor control [77].
+During sleep mode, PAC offers a proxy for modeling abstract thought (internal loop) in the absence of external stimuli; this internal loop can be extended to an external loop for social communication, which might be related to the ecological origin of natural languages [62].
 Overall, the main contributions of this paper can be summarized as follows:
-  • A geometric insight about the blessing and curse of dimensionality leads to a new understanding of specification and generalization in ML, which nicely fits the emerging direct-fit-to-nature paradigm.
-  • A construction of manifold tangling-untangling cycle (TUC) for efficiency learning.
+  • $\bullet$
+    A geometric insight about the blessing and curse of dimensionality leads to a new understanding of inductive/design bias in ML - not all overfitting is bad, some are useful.
+  • $\bullet$
+    A construction of disentanglement-entanglement cycle (DEC) for manifold learning.
 Our construction represents a novel direct-fit-to-nature approach to unsupervised learning.
 Tangling and untangling operators are implemented by context dependency and integral transform, respectively.
-  • A biological connection of TUC with sleep-wake cycle (SWC) and its related cognitive functions.
-Untangling transforms CIR to CDR for specification in the wake mode; and untangling maps CDR back to CIR for generalization in the sleep mode.
-  • An application of TUC theory to model sensorimotor and social interactions.
+  • $\bullet$
+    A biological connection of DEC with sleep-wake cycle (SWC) and its related cognitive functions.
+Untangling transforms CIR to CDR for specification in the wake mode; and untangling maps CDR back to CIR for generalization in both wake (as top-down feedback) and sleep (as memory consolidation) mode.
+  • $\bullet$
+    An application of DEC theory to model sensorimotor and hippocampal-neocortical interactions.
 We use biological findings in various studies of cognitive functions as experimental support for the new theoretical framework.
-II Geometric Intuition
-Figure 1: Kernel trick achieves linear separability by mapping input data to a higher-dimensional feature space.
-$\phi:(x,y)\rightarrow(x,y,c)$, where $c=0,1$ denote red and blue classes, is a simple lifting idea that can be generalizable to a data manifold in arbitrary dimensions.
-In graph theory, there is a well-known educational puzzle named “Three Utilities Problem”.
+2 Geometric Intuition
+Graph theory has a well-known educational puzzle named “Three Utilities Problem".
 This puzzle can be formalized as a problem in topological graph theory by asking whether the complete bipartite graph $K_{3,3}$, has a graph embedding in the plane.
 The impossibility of the puzzle corresponds to the fact that $K_{3,3}$ is not a planar graph.
-However, such a puzzle can be easily solved if the planarity constraint is removed.
-This is a simple example illustrating the blessing of dimensionality to the problem of manifold untangling.
+However, removing the planarity constraint can easily solve such a puzzle.
+This is a simple example illustrating the blessing of dimensionality to the problem of manifold disentanglement/untangling.
+In our daily lives, the overpass at the intersection is another example of entanglement/tangling - the traffic jam problem can be effectively solved by lifting the space along the vertical dimension.
+By analogy, we can imagine that the pattern classification problem can be facilitated by a similar strategy of dimensionality increase (i.e., entanglement).
+We observe that such inductive bias can be exploited to discriminate patterns in a higher dimensional space by inducing motion as the contextual variable.
+a)                                                                    b)
+Figure 2: Comparison of different geometric intuitions: a) linear separability by kernel trick. b) linear separability by hyperplane separation theorem.
 In topology, Whitney Embedding Theorem [48] states that any smooth manifold can be embedded into a Euclidean space of double its dimension.
-Based on this classical result, we can prove the following lemma:
-Lemma (Linear separability in high-dimensional space)
-Given two tangled manifolds $A,B\in R^{n}$, one can always linearly separate them in a high-dimensional space.
-Proof. Using Whitney’s embedding theorem, we can also embed $A,B$ to a Euclidean space in $\tilde{A},\tilde{B}\in R^{2n}$. Then it is trivial to linearly separate $\tilde{A},\tilde{B}$ in $R^{4n}$.
+This fundamental result implies that an increase in dimensionality can facilitate the solution to pattern classification by linearizing the representation after embedding.
+We note that the geometric intuition of embedding differs from that of kernel trick (e.g., in support vector machine [73]), as shown in Fig.
+2.
+Unlike kernel trick, embedding is independent of data representations; both can make nonlinear separability tractable by linear methods but only embedding can deliver data-agnostic solutions.
+In particular, we are interested in the class of linear dichotomies [12] as potential low-dimensional classifiers for learning task-state representations [54] after embedding.
+Based on Cover’s classical result [12], we can prove the following lemma related to the blessing of dimensionality:
+Lemma 1 (Linear separability lemma)
+Given two entangled manifolds $A,B\in R^{n}$, one can always linearly separate them in a high-dimensional space.
+Proof.
+In the hyperplane separation theorem [14], if both these sets are closed and at least one of them is compact, then there is a hyperplane in between them and even two parallel hyperplanes in between them separated by a gap (refer to Fig.
+2).
+Using Whitney’s embedding theorem, we can also embed $A,B$ to a Euclidean space in $\tilde{A},\tilde{B}\in R^{2n}$.
+Then it is trivial to linearly separate $\tilde{A},\tilde{B}$ in $R^{4n+1}$ by constructing a hyperplane along the new dimension (e.g., consider the convex hulls of $\tilde{A},\tilde{B}$).
+a)                                                         b)
+Figure 3: Abstraction of SWC by DEC and its relationship to context dependency.
+a) The disentanglement-entanglement cycle (DEC) is used to solve overfitting/underfitting problems in manifold learning, whose biological plausibility is supported by the sleep-wake cycle (SWC).
+b) Implementation of DEC by a hysteresis loop (stable states “0” and “1” correspond to CIR and CDR, respectively).
+Note that trajectories “A” and “B” uniquely determine the context dependency (i.e., symmetry-breaking) between stimulus (S) and response (R).
 In machine learning (ML), the kernel trick used by support vector machine [65] represents a good example for illustrating the blessing of dimensionality.
-As shown in Fig.
-1, linearly non-separable data sets can become untangled (linearly separable) after being mapped to a high dimensional feature space.
-The construction or learning of features has been at the foundation of all existing models, including both biological and artificial neural networks.
-Recent direct-fit models suggest that we can solve the manifold untangling problem without feature engineering [20].
-This view is also consistent with intelligence without representation [3] and supported by the discovery of concept cells [38, 37].
-As summarized in [9], identity-related feature information of an object manifold can be simultaneously conveyed by the gnostic units in the IT cortex along with contextual information (e.g., position, pose, and scale), implying a high-dimensional latent space.
-III Context Dependency and Integral Transform for Untangling and Tangling
-Our intuition is based on observing the duality between specification (overfitting) and generalization (underfitting) in learning.
-An alternative approach to over/under-fitting is direct-fit-to-nature [20], which greatly inspired our approach in this work.
-Geometrically, we interpret overfitting/underfitting as manifold untangling and tangling, as shown in Fig.
-2.
-Assuming sensory raw inputs are data belonging to a low-dimensional manifold embedded in a high-dimensional space, the objective of manifold untangling is to overfit the data for specification.
+Based on Mercer’s theorem, linearly non-separable data sets can become untangled (linearly separable) after being mapped to a high-dimensional feature space.
+The construction or learning of features has been at the foundation of many existing learning models, including biological and artificial neural networks.
+Recent direct-fit-to-nature strategy [30] suggests that we can solve the manifold untangling problem without feature engineering, such as the construction of kernel machines [73].
+The key idea is to assume a dense sampling of data manifold so generalization can be achieved by local interpolation instead of extrapolation on over-parameterized models.
+This view is also consistent with intelligence without representation [5] (i.e., “world as its own model”) and supported by the discovery of concept cells [59, 58].
+As summarized in [16], identity-related feature information of an object manifold can be simultaneously conveyed by the gnostic units or neural population geometry [11] in the IT cortex along with contextual information (e.g., position, pose, and scale), implying a high-dimensional latent space in which all decisions can be made by a collection of low-dimensional linear dichotomies [12].
+3 Entanglement and Disentanglement for Wake and Sleep
+Our biological intuition is based on observing the alternating default and perturbation modes in SWC.
+It has been hypothesized in [9] that “sleep may serve to cement or consolidate memories by replaying parts and details of the waking experience”.
+The process of memory consolidation is believed to occur during REM sleep [74], which transforms episodic memory into semantic memory.
+Episodic memories are context-dependent representations (CDR) in which context is specified by the spatiotemporal position associated with the animal’s behavior.
+Semantic memories are context-independent representation (CIR) as the consequence of generalization.
+Geometrically, $CIR\rightarrow CDR$ corresponds to dimensionality expansion because context is induced as design bias [25]; $CDR\rightarrow CIR$ is implemented by dimensionality reduction because aliasing or collapse of contextual variables leads to generalization [46].
+Computationally, we abstract the pair of dimensionality expansion and reduction operators into a cycle between CIR and CDR, as shown in Fig.
+3a.
+Assuming sensory raw inputs are data belonging to a low-dimensional manifold embedded in a high-dimensional space, the objective of manifold entanglement or tangling is to overfit the data for specification by introducing context as design/inductive bias [3].
 This objective can be facilitated by embedding the manifold into a high-dimensional space, as Whitney’s embedding theorem suggests.
-Therefore, manifold untangling maps sensory raw input to a high-dimensional latent space where the fitting task becomes trivial (e.g., by linear dichotomy [6]).
-The dual process of manifold tangling is to underfit the data for generalization.
-A common example is aliasing [29], which reduces the dimensionality by collapsing the variables with a common state.
-We suggest an integral transform as a more general strategy than aliasing for generalization.
-In this section, we will study the pair of manifold untangling and tangling separately.
-Figure 2: The tangling-untangling cycle (TUC) for solving overfitting/underfitting problems in efficient ML. Note its biological plausibility is consistent with sleep-wake cycle (SWC).
-III-A Manifold Untangling: Dimensionality Expansion via Context Dependency
-The intuition behind context dependency can be easily explained by the toy example of binary classification (e.g., cats vs.
+Therefore, manifold entanglement/tangling maps sensory raw input to a high-dimensional latent space where the task of separating different objects can be facilitated by linearization (e.g., by linear dichotomy [12]).
+The dual process of manifold disentanglement/untangling is to reduce the dimensionality of data representation for generalization.
+A common example is aliasing [46], which collapses the variables with a common state (context).
+This section will study the pair of manifold disentanglement and entanglement separately.
+3.1 Wake Mode for Specification: Dimensionality Expansion via Context Dependency
+The intuition behind context dependency can be easily explained by the toy example of binary classification in supervised learning (e.g., cats vs.
 dogs).
 Suppose we collect sensory raw data associated with two classes $\textbf{X}_{0},\textbf{X}_{1}\in R^{n}$ corresponding to classes 0 and 1, respectively.
-The raw data in $R^{n}$ is tangled (e.g., blue vs.
-red dots as shown in Fig.
-1) but they can be easily untangled by lifting the dimensionality by one - i.e., $(\textbf{X}_{c},c)\in R^{n+1}$ where $c=0,1$ (a linear classifier $c=1/2$ is sufficient).
-Such an idea of dimensionality expansion can be better articulated as a strategy of context dependency where the label is explicitly coupled with the training data as a context variable.
-Note that the introduction of label-based context dependency transforms supervised learning to unsupervised learning [2] by direct-fit-to-nature [20].
-We can generalize the strategy of context dependency for high-dimensional labels/context as follows.
-Proposition 1 (Specification). Context Dependency for Manifold Untangling.
-Given sensory raw data $\textbf{X}\in R^{n}$ and its associated label $\textbf{C}\in R^{k}$, we can untangle the tangled data distribution $P(\textbf{X})$ by introducing context dependency as a joint PDF $P(\textbf{X},\textbf{C})\in R^{n+k}$.
+The raw data in $R^{n}$ is generally not linearly separable but they can be made linearly separable by lifting the data distribution to a new dimension - i.e., $(\textbf{X}_{c},c)\in R^{n+1}$ where $c=0,1$ (a linear classifier $c=1/2$ is sufficient).
+Such an idea of dimensionality expansion can be better articulated as a strategy of context dependency that transforms a supervised learning problem into an unsupervised learning problem [2].
+The key idea is that the class labels associated with different objects can be explicitly introduced as inductive bias or context variables.
+Alternatively, we note that the spatio-temporal locality of the object often enforces the context information in the physical world.
+In other words, the local receptive field of a mammalian brain can only pay attention to a single object in space at any time.
+This strategy of overfitting with design bias implements recently proposed “direct-fit-to-nature” [30], which relies on over-parameterized models to provide mindless generalization.
+The shared vision between our context dependency and direct-fit is to densely sample the problem space so that local heuristics (e.g., nearest neighbor or averaging) will be sufficient for situating novel observations within the context of past observations.
+This vision is also consistent with intelligence without representation [5] which advocates for learning without explicitly representing underlying rules and principles in a human-interpretable way [30].
+We can generalize the strategy of direct-fit by context-dependency for high-dimensional labels/context as follows.
+Proposition 1 (Specification). Context Dependency for Direct-Fit.
+Given sensory raw data $\textbf{X}\in R^{n}$ and its associated label $\textbf{C}\in R^{k}$, we can directly fit the data distribution $P(\textbf{X})$ by introducing context dependency as a joint PDF $P(\textbf{X},\textbf{C})\in R^{n+k}$.
 The proof of Proposition 1 is a straightforward extension of the strategy used in binary classification.
 By considering the discrete representation of label or context C, linear separability is guaranteed in the expanded higher-dimensional space.
+Note that when the dimensionality of C is higher than that of X, we will observe high-dimensional encoding space as envisioned in direct-fit [30], which turns an extrapolation problem into an interpolation problem.
 […]
-III-B Manifold Tangling for Generalization: Dimensionality Reduction via Integral Transform
-The dual operation of manifold tangling can be understood as the extension of aliasing [29].
-The basic idea behind aliasing is to provide useful invariants by representing as equivalent states in which the same action/class is needed.
-Aliasing is good for generalization but has the risk of over-generalization.
-Such risk can be well managed by applying aliasing or its extended tangling operator to the untangled data in the high-dimensional space.
-The trick is to selectively remove or collapse hidden states where the context dependency dictates the criterion of selectivity.
+3.2 Sleep Mode for Generalization: Dimensionality Reduction via Integral Transform
+The dual operation of sleep mode can be understood as the extension of aliasing [46] by memory consolidation.
+The basic idea behind aliasing is to provide useful invariants or generalizations by representing equivalent states in which the same action/class is needed.
+Aliasing is usually achieved by dimensionality reduction, which collapses the distribution with shared context variables.
+More generally, we suggest that generalization can be achieved by mapping from CDR to CIR using integral transforms.
+The trick is to selectively remove or collapse hidden states C where the context dependency dictates the criterion of selectivity.
 Using contextual information C as the guardrail, we are guaranteed a faithful generalization without the risk of over-generalization.
-Proposition 2 (Generalization). Integral Transform for Manifold Tangling.
+Formally, we have the following results.
+Proposition 2 (Generalization). Integral Transform for Interpolation-based Generalization.
 […]
-The pairing of manifold untangling and tangling operators forms a complete (TUC) cycle, as shown in Fig.
-2.
-The connection of TUC with direct-fit models [20] is appealing - the objective of learning is not human-interpretable rules or world representations (i.e., “intelligence without representation” [3]).
-Instead, evolution has invented a localized computation for embodied cognition to interpolate over task-relevant manifolds (we will elaborate on this issue later when studying sensorimotor interaction).
-Manifold untangling offers a constructive solution to obtain over-parameterized models that are simple and parsimonious.
-More importantly, the connection of TUC with the evolutionary process is so powerful that we strongly believe it offers an important clue to unlock hidden secrets in developmental and ecological psychology.
-In summary, TUC forms the building block for our solution to efficient learning problems in general.
-Before we study the application of this new theory into neuroscience, we report several hierarchical extensions of TUC into chains and fundamental group of torus [21] to enrich our weaponry of constructing over-parameterized direct-fit models.
-IV Hierarchical Extension and Connection with Fractal Geometry
-First, we can spatially extend TUC with Cartesian products to form a fundamental group of the circles [21].
+Note that integral transforms are intrinsically interpolation or averaging in high-dimensional space.
+From the perspective of distributed representation, [33], we can interpret context C as a distributed set of representing X.
+In distributed representations, interpolation-based generalization happens when a neural network learns to map inputs to a high-dimensional space where similar inputs are closer together and dissimilar inputs are farther apart.
+Integral transform offers one step beyond interpolation-based generalization [33] because all noise and interference in task-irrelevant dimensions will be filtered out for improved robustness.
+The pairing of specification and generalization operators forms a complete model for SWC, as shown in Fig.
+3a.
+The connection of SWC with direct-fit models [30] is appealing - the objective of learning is not human-interpretable rules or world representations but rational behaviors for survival and prosperity in the physical world.
+Toward this objective, evolution has invented a localized computation for embodied cognition to interpolate over task-relevant manifolds (we will elaborate on this issue later when studying sensorimotor interaction).
+Direct-fit implemenyed by SWC offers a constructive solution to obtain over-parameterized models that are simple and parsimonious.
+More importantly, the connection of SWC with the evolutionary process is so powerful that we strongly believe it offers an important clue to unlock hidden secrets in developmental and ecological psychology.
+In summary, SWC forms the building block for our solution to efficient learning problems in general.
+Before we study the application of this new theory into neuroscience, we report hierarchical extensions of SWC into chains and fundamental group of torus [31] to enrich our weaponry of constructing over-parameterized direct-fit models.
+4 Hierarchical Extension: from Cycles to Torus and Chain
+First, we can spatially extend SWC with Cartesian products to form a fundamental group of the circles [31].
 Such a dimensional increase is similar to the product PDF from the two marginals.
@@
 a)                                                      b)
-Figure 3: Hierarchical extension of TUC: a) spatial extension by Cartesian product of TUC, which forms the fundamental group of the circles; b) temporal extension by alternating between untangling and tangling in time domain.
-Proposition 3. (Cartesian Product of TUC forms the Fundamental Group of the Circle).
+Figure 4: Hierarchical extension of SWC: a) spatial extension by Cartesian product of SWC, which forms the fundamental group of the circles; b) temporal extension by alternating between untangling and tangling in time domain.
+Proposition 3. (Cartesian Product of DEC forms the Fundamental Group of the Circle).
 […]
-By recursively applying Cartesian products of TUC, we can construct a topological structure with increasing dimensionalities.
+By recursively applying Cartesian products of DEC, we can construct a topological structure with increasing dimensionalities.
 Fig.
-3a illustrates a toy example for such an extension of increasing dimensionalities (from a pair of circles to a torus).
-Analogous to the extension of circles to the fundamental group of circles, we can construct the fundamental group of TUC as $\textbf{S}^{1}\times...\times\textbf{S}^{1}$.
-An alternative hierarchical extension of TUC is based on the temporal depth [35] where the process of untangling and tangling occurs in an alternating manner.
-Temporal depth or hierarchy is formed based on allostasis [35], as shown in Fig 3b.
-This construction will be useful for characterizing the dynamics of TUC, such as error feedback and temporal stability, especially in the presence of heterogeneous interactions (e.g., sensory interactions with motor).
+4a illustrates a toy example for such an extension of increasing dimensionalities (from a pair of circles to a torus).
+Analogous to the extension of circles to the fundamental group of circles, we can construct the fundamental group of DEC as $\textbf{S}^{1}\times...\times\textbf{S}^{1}$.
+An alternative hierarchical extension of DEC is based on the temporal depth [56] where the process of untangling and tangling occurs in an alternating manner.
+Temporal depth or hierarchy is formed based on allostasis [56], as shown in Fig.
+4b.
+This construction will be useful for characterizing the dynamics of DEC, such as error feedback and temporal stability, especially in the presence of heterogeneous interactions (e.g., sensory interactions with motor).
 The alternating between tangling and untangling operators is also analogous to the alternating between sleep and wake modes in sleep-wake cycles, as we will elaborate in the next section.
-Proposition 4. (Allostatic concatenation of TUCs forms temporal hierarchy.
-Temporally hierarchical extension of TUCs can support continual learning by alternating between the processes of untangling and tangling in the time domain.
-A more subtle construction of hierarchical TUC is based on fractal geometry [28].
-We must extend the strategy context dependency across different scales to support fractal TUC.
-This can be implemented by the positioning trick [26], which was inspired by Hippocampal indexing theory [46].
+Proposition 4. (Allostatic concatenation of DECs forms temporal hierarchy.
+Temporally hierarchical extension of DECs can support continual learning by alternating between the processes of untangling and tangling in the time domain.
+A more subtle construction of hierarchical DEC is based on fractal geometry [44].
+We must extend the strategy context dependency across different scales to support fractal DEC.
+The positioning trick can implement this [43], which was inspired by Hippocampal indexing theory [70].
 The indexing variables $I$ induce inter-scale context dependency - e.g., in structure-mapping theory for analogy [26], higher-order relations are modeled based on the relation between objects instead of attributes of objects (so-called systematicity principle).
 The implementation of such principle requires the context to be the object X instead of class label C, which requires a fractal design such as the six-layered organization of cortical column.
-Such hierarchical organization of structures into TUCs might explain the magical number of six (e.g., six degrees of separation) or seven [30].
-Figure 4: Illustration of sleep-wake cycle (SWC): untangling occurs during the wake time (perturbation state) and tangling occurs during the sleep time (default mode).
-V Biological Connections with PNG and SWC
-The biological implementation of the untangling operator is based on the connection of the polychronization neural groups (PNG) theory in [24] with the locally subspace untangling hypothesis in [9].
-A direct consequence of direct-fit-to-nature is over-parameterization [20] to increase the predictive power without explicitly modeling the underlying generative structure of the world (i.e., intelligence without representation [3]).
+Such hierarchical organization of structures into DECs might explain the magical number of six (e.g., six degrees of separation) or seven [47].
+5 Biological Implementations and Applications
+The biological implementation of both entanglment and disentanglement operators is based on the polychronization neural groups (PNG) theory [36].
+A direct consequence of direct-fit-to-nature is over-parameterization [30] to increase the predictive power without explicitly modeling the underlying generative structure of the world.
 PNG offers a plausible computational model for supporting such mindless yet powerful generalizations.
 The number of PNG [36] can exponentially increase as more neurons are involved.
-Intuitively, one can argue that a high-level PNG with more neurons will have more representation power than a low-level PNG, which makes it a promising candidate for implementing the direct-fit-to-nature model [20].
+Intuitively, one can argue that a high-level PNG with more neurons will have more representation power than a low-level PNG, which makes it a promising candidate for implementing the over-parameterized model required by direc-fit [30].
 We first extend Hebbian postulate [32] to make it work for PNG as follows - “A low-level PNG that $drives$ a high-level PNG to fire tends to wire together” [19].
@@
 PNG-based implementation can handle this combinatorial explosion because of its enormous capacity [36].
-[…]
-a)                                                      b)
-Figure 5: Context dependency matters to both visual perception and motor control. a) PNG-based untangling for visual perception; b) PNG-based untangling for motor control.
-The combination of PNG with SWC [13] allows us to study cognitive functions from a novel TUC perspective.
-Using TUC as the vocabulary, we attempt to construct a computational model for SWC as the first step toward a more comprehensive neural syntax [4].
+The connection of PNG with SWC [21] allows us to study cognitive functions from a novel DEC perspective.
+Using DEC as the vocabulary, we attempt to construct a computational model for SWC as the first step toward a more comprehensive neural syntax [9].
 As application examples, we consider three tasks here: spatial navigation, object recognition, and motor control.
@@
 Untangling during the wake mode (perturbation state) is demonstrated by place cells with unidirectional place fields.
-The 1D travel sequence dictates the spatio-temporal context, which helps resolve the ambiguity in space [4].
-Tangling during the sleep mode (default state) is demonstrated by omnidirectional place cells [34] that do not require a temporal context.
-Memory consolidation [44, 49] implements the aliasing or tangling by binding all PNG which shared the same context C together.
+The 1D travel sequence dictates the spatio-temporal context C, which helps resolve the ambiguity in space [9].
+Tangling during the sleep mode (default state) is demonstrated by omnidirectional place cells X [55] that do not require a temporal context C.
+Memory consolidation [69, 74] implements the aliasing or tangling by binding all PNG which share the same context C together.
 The collapse of cloned variables occurs at the hippocampus where the encoding of spatio-temporal context (index to the local map or position on the global map) is handled by CA3-CA1 collateral system (pp.
@@
 Empirical evidence supporting this line of theoretical reasoning can be found in [13].
-VI Computational Modeling of Sensorimotor and Social Interactions
-TUC-based local computation for direct-fit inspired us to propose a new conceptual framework called “localized embodied cognition” with the locality principle in mind from problem formulation to solution algorithms.
+a)                                                      b)
+Figure 5: Context dependency matters to both visual perception and motor control. a) PNG-based untangling for visual perception; b) PNG-based untangling for motor control.
+6 Computational Modeling of Sensorimotor and Social Interactions
+DEC-based local computation for direct-fit inspired us to propose a new conceptual framework called “localized embodied cognition” with the locality principle in mind from problem formulation to solution algorithms.
 We still use the world “as its own model” [5], but the way local embodied cognition never attempts to solve a global problem locally.
@@
 By connecting with different sensory and motor areas, they offer a universal processing algorithm [53].
-TUC-based modeling allows us to gain a deeper understanding of sensorimotor interaction at the system level without the distraction from lower-level implementations.
+DEC-based modeling allows us to gain a deeper understanding of sensorimotor interaction at the system level without the distraction from lower-level implementations.
 As shown in Fig.
@@
 Conjecture (Context-dependent manifold untangling) The task of sensorimotor learning can be facilitated by simultaneously untangling the neural manifold of sensory and motor variables in the right positioning of latent space, where the position is specified by the index from the hippocampus as in Hippocampal indexing theory [70].
-SMC-based Social Interaction
-.
+SMC-based Social Interaction.
 Carrying the “more is different” principle one step further, we can envision a bigger cycle connecting a pair of SMCs.
@@
 Even though within each neocortex, the hierarchical organization maximizes the information flow inward, social communication (flowing outward) has to wait for evolution to improve its efficiency.
-Figure 7: The cycle connecting speech production by motor cortex and language comprehension by auditory cortex has a dual structure to SMC because social communication between two people is connected by CIR (tangled states).
 The fundamental role played by natural language is to facilitate communication among people - from non-verbal to verbal.
-Fig.
-7 shows a modification of SMC to abstract the verbal communication between two people A and B, where motor and auditory cortex are responsible for speech generation and language comprehension, respectively.
-Note that unlike the original SMC in Fig.
-6, the interface between two TUCs is in the tangled state (CIR).
-This difference has a fundamental impact on our internal (self-conscious) and external (social) cognitive function.
-For internal communication between sensory and motor cortex, the interface (inverse and forward in Fig.
-6) connects two untangled states and there is no bottleneck in terms of information flow.
-For external (social) communication between two different person, the interface (inverse and forward in Fig.
-7) connects two tangled states and faces an information flow bottleneck.
-According to the simulation theory of mind-reading [15], mirror neurons can help us understand and empathize with the emotions and intentions behind others’ non-verbal cues.
-Modeling the new cycle in Fig.
-7for understanding the social nature of language might offer new insight to the neural exploitation hypothesis [14].
-According to this hypothesis, social cognition and language are both connected with action through the support of mirror neurons.
-However, the difference between non-verbal and verbal communication seems to be determined by their organizational principles.
-Non-verbal communication such as sign language does not admit hierarchical organization; while verbal communication such as Chomsky’s hierarchy [5] in natural language offers an evolutionary solution to maximize the information flow within the new cycle.
-VII Conclusions and Future Works
-In this paper, we have constructed a new TUC-based theory to solve the long-standing learning problem.
-Exploiting the blessing of dimensionality by inducing context dependency allows us to untangle any data manifold so linear classifiers will be sufficient.
-Then the tangling operator based on integral transform collapses the induced context variables for generalization.
-Such mapping between tangled CIR and untangled CDR can be extended into hierarchical representations by Cartesian product and fractal geometry.
-The computational efficiency of TUC-based learning algorithms is supported by the biological implementation by SWC.
-Sleep and wake modes in SWC elegantly correspond to tangling and untangling operators due to the absence of sensory inputs during the sleep mode.
-SWC allows us to construct a neural syntax for an improved understanding of cognitive functions associated with sensorimotor interaction and social interaction.
+If one concatenates the SMCs of two people (for communication purposes), we observe a larger cycle $SMC_{A}\leftrightarrows SMC_{B}$.
+According to the simulation theory of mind-reading [24], mirror neurons can help us understand and empathize with the emotions and intentions behind others’ nonverbal cues.
+Modeling the bigger cycle (refer to Appendix A for a more detailed discussion about the ecological origin of languages) for understanding the social nature of language might offer new insight to the neural exploitation hypothesis [23].
+Since nature is good at recycling its tricks, the originally evolved sensorimotor integration as shown in Fig.
+6 might be adapted for the new function of language as a more efficient alternative to nonverbal communication.
+7 Conclusions, Limitations, and Broader Impact
+Conclusions: This paper studies the computational modeling of sleep-wake cycle (SWC) from a direct-fit perspective.
+The wake mode entangles the sensory observations by binding them into gnostic cells.
+The sleep mode may consolidate memories by replaying the waking experience.
+Memory consolidation is abstracted by an integral transform in this paper, which disentangles the manifold for generalization.
+The pair of entanglement and disentanglement operators implement the mapping between CIR and CDR, demonstrating the importance of context dependency to learning and memory.
+The hierarchical extension of disentanglement-entanglement cycles into high-order topological structures (e.g., torus and chain) is an attempt to enhance the computational modeling capability of SWC for more sophisticated cognitive tasks.
+Limitations: Our paper is inspired by many cycles discovered by Buzsaki in the rhythms of the brain [9].
+Its biggest limitation is the lack of experimental verification.
+As pointed out in [9], studying the brain’s default mode is even more challenging than the perturbation mode.
+Experiments related to sleep-dependent memory consolidation [69] only indirectly provide evidence to support the role of SWC in semantic memory.
+Many cycles discovered in this paper are limited to theoretical appeal, whose practical implications and connections remain to be established.
+Due to space limitation, we have skipped the connection of DEC with limit-cycle behavior of Wilson-Cowan model [76] and its related hysteresis cycle (the physical origin of memory).
+Broader Impacts: Understanding the sleep-wake cycle paves the path to unlocking the secret of learning and memory.
+At the device level of individual neurons, limit-cycle behavior has been well studied by the Wilson-Cowan model [76].
+At the circuit level of neural groups, canonical circuits [39] have formed the cycle for predictive processing.
+This paper marks the first step toward discovering cycles at the system level of neural areas.
+Since SWC is fundamental to almost all cognitive functions, its computational modeling might facilitate the study of both normal behaviors and neurological disorders such as autism and anxiety.
+Studying the default state of the brain (sleep mode) is deemed to complement the study of the perturbation state (wake mode).
+Appendix A Ecological View of Sleep/Wake Modes and Sleep-Wake Cycle
+The Brain’s Default State: Sleep Mode.
+According to Gyorgy Buzsaki [9], rest and sleep are characterized by self-organized oscillations in the absence of environmental inputs.
+Without an outside influence, the evolving state in the sleep mode is deterministic (i.e., completely predictable).
+The role of sleep in memory consolidation has been observed in the experiments with sleep spindles [64].
+It was hypothesized that during spindles, specific neuronal populations become more strongly modulated through virtual replay [37].
+Experiments with the visual cortex and hippocampus have suggested that they support simultaneous reactivation of coherent memory traces during slow-wave sleep (SWS) and rapid-eye-movement (REM) sleep [37].
+All these experimental findings are consistent with Proposition 2 in our SWC theory.
+In summary, Memory consolidation is a process by which the hippocampus guides the organization of the information stored in the neocortex [68].
+The consequence of self-organization is that the consolidated memory becomes independent of the hippocampus - e.g., from episodic memory (context-dependent) to semantic memory (context-independent).
+The Brain’s Perturbation State: Wake Mode.
+According to Karl Friston, “brain dynamics constantly shift from the complex to the predictable” (pp.
+111, [9]).
+The waking brain has been hypothesized to be optimized for encoding memories [60].
+From sensory perception to motor control, goal-oriented behaviors offer a unifying perspective of understanding the brain’s state during the wake mode.
+According to Hebbian postulate [32], the synchronization of cell assemblies (a.k.a., polychronization of neural groups [36]) drives the brain dynamics “from stable incoherence through dynamic instability to complete entrainment” [20].
+Unlike the deterministic default mode, the stochastic nature of environmental inputs dictates the brain to strike a good balance between specification and generalization.
+By encoding history into memory, complex systems manage to make reliable predictions.
+Proposition 1 in our SWC theory is consistent with this line of reasoning because CDR facilitates the prediction by inductive bias.
+Sleep-Wake Cycle and Memory Consolidation.
+Not only humans but also mammals go to sleep [66].
+Across mammals, the amount and nature of sleep are often correlated with their body size, age, and ecological variables (e.g., terrestrial or aquatic).
+The role of the SWC in energy conservation and nervous system recuperation has been widely observed in mammals.
+The recognition of SWC as a biological rhythm controlled by brain oscillators (i.e., beyond the framework of homeostatic principles) has only been more widely accepted since the 1960s [42].
+The basis of SWC has been related to the process of toxins’ accumulation and removal [42].
+The neurobiology of SWC was reviewed in [21], highlighting the sleep architecture, circadian regulation, and regulatory feedback.
+The close relationship between memory encoding/consolidation and SWC was studied by Buzsaki in 1998 [7] (a more recent review can be found in [17]).
+The integral transform in Proposition 2 might be a mathematical abstraction of synaptic modifications in the hippocampus during sleep.
+The importance of sleep patterns to the preservation of experience-induced synaptic modifications has been documented by various studied on sleep deprivation [40].
+Appendix B On Ecological Origin of Languages
+The origin of language was once a taboo among linguists due to a lack of scientific methodology [34].
+In 1950, Alan Turing invented an imitation game (later known as the Turing test) for artificial intelligence (AI) [72].
+In the Turing test, a human evaluator would judge natural language conversations between a human and a machine designed to generate human-like responses.
+Recent breakthroughs in generative AI (e.g., ChatGPT developed by OpenAI [6]) have shown convincing evidence for machine’s nearly passing the Turing test (“Essentially, whether I pass the Turing Test might depend on the context and the criteria set by the human interacting with me.” - quoted from ChatGPT).
+However, passing a Turing test does not imply the possession of minds like humans do, as demonstrated by John Seale in Chinese room thought experiments in 1980.
+The fundamental reason behind the Chinese room argument is that the Turing test is a top-down approach to understanding intelligence [9].
+An alternative approach is bottom-up (a.k.a.
+inside-out), which requires a deeper understanding of the mechanism underlying the structure and function of neural systems.
+As discovered by Vernon Moutcastle in 1957 [52], there is no visible difference among the six-layer organization of the neocortex no matter which area we look into.
+Based on this observation, he hypothesized that all cognitive functions, from spatial navigation and motor control to vision and language, result from a universal cortical processing algorithm.
+According to Theodosius Dobzhansky, “nothing in biology makes sense except in the light of evolution.” The evolution of mammalian brains offers fruitful insight into the functional organization of the neocortex.
+As Hans Moravec says in [49], abstract thought “is a new trick, perhaps less than 100 thousand years old….effective only because it is supported by this much older and much more powerful, though usually unconscious, sensorimotor knowledge.” To understand language, we must understand sensorimotor learning first.
+Following the clue offered by the evolution of mammalian brains, one can understand the allocortex that is sandwiched between the old and new brain for spatial navigation [71].
+The discovery of place cells in the hippocampus in 1978 [55] marked a milestone in our journey to a better understanding of how the brain works.
+This line of research has accumulated many more discoveries (e.g., grid cells [50, 51]) and computational models related to navigation, contributing to our knowledge about how to build cognitive maps [75].
+Along this line of reasoning, the relationship between place cells and memory consolidation [69] is particularly fascinating because if Mountcastle were correct, the trick used by nature to create gnostic units such as place cells is likely to be recycled for other cognitive functions (e.g., concept cells in object recognition [59, 58]).
+To solve the navigation problem, context (for specification) and aliasing (for generalization) are two clever tricks invented by nature.
+As shown in [9] (pp.
+327), oscillatory timing of cell assemblies defines the spatiotemporal context for a currently coded item (e.g., the current position of the animal).
+During the wake mode (perturbation state), animals generate unidirectional place fields as context-dependent episodic memories about one-dimensional travel routes.
+During the sleep mode (default state), memory consolidation uses an aliasing trick to collapse unidirectional place cells with identical positions but different contexts into omnidirectional place cells.
+Unlike unidirectional place cells that are context-dependent, omnidirectional place cells are context-independent, implying better generalization.
+Such an aliasing trick becomes Maslow’s hammer [45] to tackle other cognitive tasks.
+a)                                                         b)
+Figure A1: Ecological origin of languages.
+a) language is needed to support the external sensorimotor cycle where the sensory cortex and motor cortex cannot directly communicate with each other but have to overcome the information bottleneck of the auditory cortex (language comprehension is the inverse pass transforming auditory-CIR into motor-CIR to facilitate motor control for speech generation); b) Chomsky hierarchy mirrors the nested organization of the physical world so the neocortex can recycle Maslow’s hammer in verbal communication.
+Nature has cleverly used the aliasing trick, based on Maslow’s hammer, to solve the learning problem by memory.
+In simple terms, the spatio-temporal context in navigation becomes a response vector to a feature input in perception or action.
+The problem of sensorimotor learning is then transformed into a generalized navigation problem of memorizing a “travel sequence” in the latent space [67] that comprises the concatenation of feature input and response vectors.
+During the wake mode, the brain encodes many such travel sequences, which are known to suffer from overfitting.
+However, during sleep mode, memory consolidation collapses episodic memory into semantic memory.
+This is achieved by binding polychronization neural groups with the same feature input but different responses (contextual variables) together for generalization.
+During sleep-wake cycles [74, 69], overfitted context-dependent representations (CDR) are self-organized into context-independent representations (CIR) without the bias-variance barrier [25] or the curse of dimensionality [4] because the brain does not solve any optimization problem but count on memory (Maslow’s Hammer) for adaptive behaviors.
+The evolution of sensorimotor learning greatly contributes to human flexible behaviors in terms of perception and action [67].
+However, social communication between two people requires another clever recycling of Maslow’s Hammer.
+Unlike abstract thoughts that can be done internally in our mind, social communication requires the completion of an external cycle connecting the sensorimotor systems of two people.
+Such an external loop is at the mercy of the information bottleneck of our sensory organs because we cannot directly communicate with others like a three-body civilization (in SCIFI fiction “Three-Body”) that cannot lie.
+Instead, we have to count on external communication channels to exchange our thoughts.
+Visual communication such as sign language is limited by its capacity; natural language is more effective because the neocortex can recycle its sensorimotor skills learned from interacting with the world (corresponding to the internal cycle, as shown Fig.
+6) to interacting with the peer (corresponding to the external cycle, as shown Fig.
+A1a).
+Such ecological explanation for the origin of natural languages is another evidence supporting Maslow’s Hammer used by evolution.
+This line of reasoning explains the hierarchical organization discovered by Noam Chomsky in 1957 [10] and supports the Sapir–Whorf hypothesis [38] in linguistic relativity.
+The Sapir–Whorf hypothesis, also known as linguistic relativity, is a concept in linguistics that suggests the structure of a language affects its speakers’ worldview or cognition.
+In summary, Chomsky hierarchy (Fig.
+A1b) mirrors the nested organization of the physical world.